OpenStudy (anonymous):

What is the simplest form of the number? 32^4/5

OpenStudy (campbell_st):

32 2^5 so you have \[(2^5)^{\frac{4}{5}}\] us the index law for power of a power \[(x^a)^b = x^{a \times b}\]